OpenStudy (anonymous):

What is the correct solution to the equation? 4(x – 3) = 16 A. 1 B. 7 C. 15 D. 24

OpenStudy (anonymous):

4(x-3)=16 x-3=16/4 x-3=4 x=4+3 x=7 Answer: 7

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

7 xx

OpenStudy (anonymous):

thank you